# MARS: Unleashing the Power of Variance Reduction for Training Large Models

### v3

@@ 2 Preliminaries
$\mathbf{m}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\widetilde{\mathbf{x}},\bm{\xi}_{t})+\nabla F(\widetilde{\mathbf{x}}),$
where $\widetilde{\mathbf{x}}$ is an anchoring point (a.k.a., reference point) that updates periodically. This variance-reduced stochastic gradient can reduce the variance of the stochastic gradient by adding a correction term $-\nabla f(\widetilde{\mathbf{x}},\bm{\xi}_{t})+\nabla F(\widetilde{\mathbf{x}})$ based on a less frequently updated reference point $\widetilde{\mathbf{x}}$ and its full gradient $\nabla F(\widetilde{\mathbf{x}})$. It can be shown that the variance $\mathbf{m}_{t}$ can be controlled by $\|\mathbf{x}_{t}-\widetilde{\mathbf{x}}\|_{2}$, which will diminish as both $\mathbf{x}_{t}$ and $\widetilde{\mathbf{x}}$ converges to the stationary points when the algorithm makes progress. Subsequent improvements in variance reduction techniques were introduced in SARAH (Nguyen et al., 2017a) and SPIDER (Fang et al., 2018), which get rid of the anchor point and result in the following momentum update:
@@ 3.1 MARS Framework
$X$ $=(1-\beta)\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})$
$Y$ $=\beta\left[\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla f(\mathbf{x}_{t},\bm{\xi}_{t+1})\right]$
$Z_{t}$ $=\mathbf{m}_{t}-\nabla F(\mathbf{x}_{t})$
@@
On the other hand, Adam can be seen as using a diagonal $\mathbf{H}_{t}$, where each diagonal element is the EMA of $[\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})]_{i}^{2}$:
@@
Therefore, the update simplifies to elementwise adaptive gradient update, i.e., $[\mathbf{x}_{t+1}]_{i}=[\mathbf{x}_{t}]_{i}-\eta[\mathbf{m}_{t}]_{i}/[\mathbf{H}_{t}]_{ii}$. Our unified framework accommodates both types of preconditioning: full Hessian approximation and diagonal Hessian approximation. Different definitions of $\mathbf{H}_{t}$ give rise to different algorithms.
@@ 3.2.1 MARS-AdamW
$\widehat{\mathbf{m}}_{t}$ $=\frac{\mathbf{m}_{t}}{1-\beta_{1}^{t}},~{}~{}\widehat{\mathbf{v}}_{t}=\frac{\mathbf{v}_{t}}{1-\beta_{2}^{t}},$ (3.9)
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\eta_{t}\bigg{(}\frac{\widehat{\mathbf{m}}_{t}}{\sqrt{\widehat{\mathbf{v}}_{t}}+\epsilon}+\lambda\mathbf{x}_{t}\bigg{)}.$ (3.10)
@@
Here, $\gamma_{t}$ is a scaling parameter that controls the strength of gradient correction. When $\gamma_{t}=0$, the algorithm reduces to AdamW. Conversely, when $\gamma_{t}=1$, (3.13) aligns with the STORM momentum. Combining (3.12), (3.13), (3.14) together with (3.11) and the mirror descent update (3.2), we derive the MARS-AdamW algorithm in Algorithm 2. In practice, $\gamma_{t}$ is often set between 00 and $1$. Moreover, we employ gradient clipping-by-norm to $\mathbf{c}_{t}$ at Line 5, following the standard gradient clipping technique performed in neural network training. We provide a convergence analysis of Algorithm 2 in Theorem B.6 in Appendix B.2.
@@
5:     if $\|\mathbf{c}_{t}\|_{2}>1$, then $\tilde{\mathbf{c}}_{t}=\frac{\mathbf{c}_{t}}{||\mathbf{c}_{t}||_{2}}$ else $\widetilde{\mathbf{c}}_{t}=\mathbf{c}_{t}$
@@ 3.2.2 MARS-Lion
$\mathbf{m}_{t}$ $=\beta_{2}\mathbf{u}_{t-1}+(1-\beta_{2})\nabla f(\mathbf{x}_{t},\bm{\xi}_{t}),$ (3.15)
$\mathbf{u}_{t}$ $=\beta_{1}\mathbf{u}_{t-1}+(1-\beta_{1})\nabla f(\mathbf{x}_{t},\bm{\xi}_{t}),$ (3.16)
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\eta_{t}\Big{(}\text{sign}(\mathbf{m}_{t})+\lambda\mathbf{x}_{t}\Big{)}.$ (3.17)
@@
Setting $\mathbf{m}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$, $a_{1}=\beta_{1}$, $a_{2}=1-\beta_{1}$, $b_{1}=\beta_{2}$, $b_{2}=1-\beta_{2}$ in Lemma D.1, and shifting the index of $\mathbf{u}_{t}$ by taking $\mathbf{u}_{t}=\mathbf{u}_{t-1}$, $\mathbf{u}_{t-1}=\mathbf{u}_{t-2}$ in Lemma D.1, we can show that Lion momentum updates in (3.15) and (3.16) are equivalent to the following single momentum update:
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$ (3.19)
$\qquad+\frac{\beta_{1}(1-\beta_{2})}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})\big{)}\Big{]}.$ (3.20)
On the other hand, setting $\gamma_{t}=1-\beta_{2}$ in the core updates of MARS (3.12) and (3.13), we obtain the $\mathbf{m}_{t}$ update:
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$ (3.21)
$\qquad+\frac{\beta_{1}(1-\beta_{2})}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}\Big{]}.$ (3.22)
The only difference between (3.26) and (3.28) lies in the stochasticity used, specifically, $\bm{\xi}_{t}$ versus $\bm{\xi}_{t-1}$ when calculating $\nabla f(\bm{x}_{t-1},\cdot)$. Therefore, ignoring the gradient clipping at Line 5, we can see Lion as a special case of MARS-Lion when $\gamma_{t}=1-\beta_{2}$, and using approximate gradient calculation on $\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})$. In practice, we observe little difference between using $f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})$ derived from the STORM momentum and its approximation $f(\mathbf{x}_{t-1},\bm{\xi}_{t})$.
@@
Setting $\mathbf{g}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$, $a_{1}=\beta_{1}$, $a_{2}=1-\beta_{1}$, $b_{1}=\beta_{2}$, $b_{2}=1-\beta_{2}$ in Lemma D.1, and shifting the index of $\mathbf{u}_{t}$ by taking $\mathbf{u}_{t}=\mathbf{u}_{t-1}$, $\mathbf{u}_{t-1}=\mathbf{u}_{t-2}$ in Lemma D.1, we can show that Lion momentum updates in (3.15) and (3.16) are equivalent to the following single momentum update:
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$ (3.25)
$\qquad+\frac{\beta_{1}(1-\beta_{2})}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})\big{)}\Big{]}.$ (3.26)
On the other hand, setting $\gamma_{t}=1-\beta_{2}$ in the core updates of MARS (3.12) and (3.13), we obtain the $\mathbf{m}_{t}$ update:
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{m}_{t-1}+(1-\beta_{1})\Big{[}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$ (3.27)
$\qquad+\frac{\beta_{1}(1-\beta_{2})}{1-\beta_{1}}\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}\Big{]}.$ (3.28)
The only difference between (3.26) and (3.28) lies in the stochasticity used, specifically, $\bm{\xi}_{t}$ versus $\bm{\xi}_{t-1}$ when calculating $\nabla f(\bm{x}_{t-1},\cdot)$. Therefore, ignoring the gradient clipping at Line 5, we can see Lion as a special case of MARS-Lion when $\gamma_{t}=1-\beta_{2}$, and using approximate gradient calculation on $\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})$. In practice, we observe little difference between using $f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})$ derived from the STORM momentum and its approximation $f(\mathbf{x}_{t-1},\bm{\xi}_{t})$.
@@ B.2 Convergence of MARS
In Algorithm 2, under Assumptions B.1B.2 and B.3, when choosing $\eta_{t}=(s+t)^{-1/3},s\geq\max(8L^{3}/\rho^{3},64\lambda^{3})$. Suppose $||\mathbf{x}_{t}||_{2}\leq D$, $c\geq 32L^{2}\rho^{-2}+1$, $\beta_{1,t+1}=1-c\eta_{t}^{2}$ and $\beta_{2,t+1}=1-\eta_{t}^{6}$, then $\forall T\geq s$, it holds that
@@
The theorems above guarantee the convergence rate of $O(\log(T)/T^{1/3})$. We remark that even though it seems that the involvement of weight decay may results in slower convergence, it performs better in practice, à la Loshchilov & Hutter (2019). Finally, the term $M_{t+1}$ is always greater or equal to 00, and when $\gamma_{t+1}=1$, we would have $M_{t+1}=0$. The dependency of $M_{t+1}$ on $\gamma_{t+1}$ is shown in (C.2) and Lemma C.2. This proves that the convergence is faster if we allow a flexible $\gamma_{t}$ schedule.
@@ Appendix D Proof of Auxiliary Lemmas
D.1 Lemma D.1 and Proof
@@
For any sequence $\{\mathbf{m}_{t}\in\mathbb{R}^{d}\}_{t=0,1,\ldots}$, consider the following updates of $\mathbf{m}_{t}$ for any constant factors $a_{1},a_{2},b_{1}$, and $b_{2}$:
@@
For $I$, we observe that $\bm{\varepsilon}_{t}$ is independent of $\bm{\xi}_{t+1}$ and the expectations of $\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})$ and $\bm{\Delta}_{t}-\mathbb{E}\bm{\Delta}_{t}$ are all 00. Therefore
@@
The second inequality follows from applying both the Cauchy-Schwarz and Young’s inequalities. Moreover, the final inequality results from selecting $\eta_{t}$ to satisfy $\eta_{t}<\frac{\rho}{2L}$ This completes our proof.
@@ E.4 Computer Vision Experiments
To explore the impact of $\gamma_{t}$, we test various $\gamma$s on GPT-2 small model, including constant and linearly changing schedules. And we plot the training and validation curves in Figure 14. It can be observed that there are slight differences among different $\gamma$s where 0.025 is the best $\gamma$. Therefore, we used $\gamma=0.025$ for other experiments in this paper.
Figure 14: Training loss and validation loss curves for MARS and MARS-approx on GPT-2 small (125M) models trained with MARS-AdamW-approx with different $\gamma$s, pretrained with OpenWebText dataset and plotted against training tokens.

### v4

@@ 2 Preliminaries
[…]
@@ 3.1 MARS Framework
$X$ $=(1-\beta)\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1}),$
$Y$ $=\beta\left[\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla f(\mathbf{x}_{t},\bm{\xi}_{t+1})\right],$
$Z_{t}$ $=\mathbf{m}_{t}-\nabla F(\mathbf{x}_{t}),$
@@
On the other hand, Adam can be seen as using a diagonal $\mathbf{H}_{t}$, where each diagonal element is the EMA of $\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})]_{i}^{2}$:
@@
Therefore, the update simplifies to elementwise adaptive gradient update, i.e., $\mathbf{x}_{t+1}]_{i}=[\mathbf{x}_{t}]_{i}-\eta[\mathbf{m}_{t}]_{i}/[\mathbf{H}_{t}]_{ii}$. Our unified framework accommodates both types of preconditioning: full Hessian approximation and diagonal Hessian approximation. Different definitions of $\mathbf{H}_{t}$ give rise to different algorithms.
@@ 3.2.1 MARS-AdamW
$\widehat{\mathbf{m}}_{t}$ $=\frac{\mathbf{m}_{t}}{1-\beta_{1}^{t}},~~\widehat{\mathbf{v}}_{t}=\frac{\mathbf{v}_{t}}{1-\beta_{2}^{t}},$
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\eta_{t}\bigg{(}\frac{\widehat{\mathbf{m}}_{t}}{\sqrt{\widehat{\mathbf{v}}_{t}}+\epsilon}+\lambda\mathbf{x}_{t}\bigg{)}.$
@@
Here, $\gamma_{t}$ is a scaling parameter that controls the strength of gradient correction. When $\gamma_{t}=0$, the algorithm reduces to AdamW. Conversely, when $\gamma_{t}=1$, (3.11) aligns with the STORM momentum. Combining (3.10), (3.11), (3.12) together with (3.9) and the mirror descent update (3.2), we derive the MARS-AdamW algorithm in Algorithm 2. In practice, $\gamma_{t}$ is often set between 0 and $1$. Moreover, we employ gradient clipping-by-norm to $\mathbf{c}_{t}$ at Line 5, following the standard gradient clipping technique performed in neural network training. We provide a convergence analysis of Algorithm 2 in Theorem B.6 in Appendix B.2.
@@
[…]
@@ 3.2.2 MARS-Lion
$\mathbf{m}_{t}$ $=\beta_{1}\mathbf{u}_{t}+(1-\beta_{1})\nabla f(\mathbf{x}_{t},\bm{\xi}_{t}),$ (3.13)
$\mathbf{u}_{t+1}$ $=\beta_{2}\mathbf{u}_{t}+(1-\beta_{2})\nabla f(\mathbf{x}_{t},\bm{\xi}_{t}),$ (3.14)
$\mathbf{x}_{t+1}$ $=\mathbf{x}_{t}-\eta_{t}\Big{(}\text{sign}(\mathbf{m}_{t})+\lambda\mathbf{x}_{t}\Big{)}.$
@@
$\mathbf{m}_{t}$ $=a_{1}\mathbf{m}_{t-1}+(b_{1}a_{2}-a_{1}b_{2}+b_{2})\mathbf{g}_{t}$
$\qquad+(a_{1}b_{2}-b_{1}a_{2})(\mathbf{g}_{t}-\mathbf{g}_{t-1}).$
By setting $\mathbf{g}_{t}=\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$, $a_{1}=\beta_{2}$, $a_{2}=1-\beta_{2}$, $b_{1}=\beta_{1}$, and $b_{2}=1-\beta_{1}$ in Lemma 3.4, we can show that Lion momentum updates in (3.13) and (3.14) are equivalent to the following single momentum update:
$\mathbf{m}_{t}$ $=\beta_{2}\mathbf{m}_{t-1}+(1-\beta_{2})\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$
$\qquad+(\beta_{2}-\beta_{1})\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})\big{)}.$ (3.18)
On the other hand, by redefining $\beta_{1}=\beta_{2}$, $\beta_{2}=\beta_{1}$, and setting $\gamma_{t}=\tfrac{\beta_{2}-\beta_{1}}{\beta_{2}}$ in the core updates of MARS (3.10) and (3.11), we obtain the update for $\mathbf{m}_{t}$:
$\mathbf{m}_{t}$ $=\beta_{2}\mathbf{m}_{t-1}+(1-\beta_{2})\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})$
$\qquad+(\beta_{2}-\beta_{1})\big{(}\nabla f(\mathbf{x}_{t},\bm{\xi}_{t})-\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})\big{)}.$ (3.19)
The only difference between (3.18) and (3.19) lies in the stochasticity used, specifically, $\bm{\xi}_{t}$ versus $\bm{\xi}_{t-1}$ when calculating $\nabla f(\bm{x}_{t-1},\cdot)$. Therefore, ignoring the gradient clipping at Line 5, we can see Lion as a special case of MARS-Lion using approximate gradient calculation on $\nabla f(\mathbf{x}_{t-1},\bm{\xi}_{t})$. In practice, we observe little difference between using $f(\mathbf{x}_{t-1},\bm{\xi}_{t-1})$ derived from the STORM momentum and its approximation $f(\mathbf{x}_{t-1},\bm{\xi}_{t})$.
@@ B.2 Convergence of MARS
In Algorithm 2, under Assumptions B.1B.2 and B.3, when choosing $\eta_{t}=(s+t)^{-1/3},s\geq\max(8L^{3}/\rho^{3},64\lambda^{3})$. Suppose $\mathbf{x}_{t}||_{2}\leq D$, $c\geq 32L^{2}\rho^{-2}+1$, $\beta_{1,t+1}=1-c\eta_{t}^{2}$ and $\beta_{2,t+1}=1-\eta_{t}^{6}$, then $\forall T\geq s$, it holds that
@@
The theorems above guarantee the convergence rate of $O(\log(T)/T^{1/3})$. We remark that even though it seems that the involvement of weight decay may results in slower convergence, it performs better in practice, à la Loshchilov & Hutter (2019). Finally, the term $M_{t+1}$ is always greater or equal to 0, and when $\gamma_{t+1}=1$, we would have $M_{t+1}=0$. The dependency of $M_{t+1}$ on $\gamma_{t+1}$ is shown in (C.2) and Lemma C.2. This proves that the convergence is faster if we allow a flexible $\gamma_{t}$ schedule.
@@ Appendix D Proof of Auxiliary Lemmas
D.1 Proof of Lemma 3.4
Proof of Lemma 3.4.
Shifting the index of (3.17) by one and bringing into (3.16), we obtain
$\mathbf{m}_{t}=b_{1}\big{(}a_{1}\mathbf{u}_{t-1}+a_{2}\mathbf{g}_{t-1}\big{)}+b_{2}\mathbf{g}_{t}=a_{1}b_{1}\mathbf{u}_{t-1}+b_{2}\mathbf{g}_{t}+b_{1}a_{2}\mathbf{g}_{t-1}.$ (D.1)
On the other hand, shifting the index of (3.16) by $1$, it holds that
$\mathbf{m}_{t-1}=b_{1}\mathbf{u}_{t-1}+b_{2}\mathbf{g}_{t-1}.$ (D.2)
$\mathbf{m}_{t}$ $=a_{1}\mathbf{m}_{t-1}-a_{1}b_{2}\mathbf{g}_{t-1}+b_{2}\mathbf{g}_{t}+b_{1}a_{2}\mathbf{g}_{t-1}$
$=a_{1}\mathbf{m}_{t-1}+(b_{1}a_{2}-a_{1}b_{2}+b_{2})\mathbf{g}_{t}+(a_{1}b_{2}-b_{1}a_{2})(\mathbf{g}_{t}-\mathbf{g}_{t-1}).$
This completes the proof. ∎
D.2 Lemma D.1 and Proof
@@
For $I$, we observe that $\bm{\varepsilon}_{t}$ is independent of $\bm{\xi}_{t+1}$ and the expectations of $\nabla f(\mathbf{x}_{t+1},\bm{\xi}_{t+1})-\nabla F(\mathbf{x}_{t+1})$ and $\bm{\Delta}_{t}-\mathbb{E}\bm{\Delta}_{t}$ are all 0. Therefore
@@
The second inequality follows from applying both the Cauchy-Schwarz and Young’s inequalities. Moreover, the final inequality results from selecting $\eta_{t}$ to satisfy $\eta_{t}<\frac{\rho}{2L}$ . This completes our proof.
@@ E.4 Computer Vision Experiments
[…]